# spacing=nonfrench

### v1

@@ VI-A Qualitative Insights
Bug Scenario GPT-4 Chronos Chronos Resolution Path
@@
TABLE XII: Qualitative examples where Chronos successfully applies multi-code context to resolve debugging tasks beyond the reach of baseline LLMs.
VI-B Ablation Studies
@@
FullNo MCAStaticNo Orch002020202040404040606060608080808010010010010090909090494949496262626255555555ConfigurationDebug Success (%)
@@
VI-C Failure Analysis
@@
VI-C1 Common Failure Modes
@@
VI-C2 Edge Cases and Limitations
  • Extremely Large Monorepos (¿10M LOC): Performance degrades to 45.3% success rate due to retrieval precision issues
@@ VI-C2 Edge Cases and Limitations
TABLE XIII: Chronos performance on challenging bug categories.
VII LIMITATIONS, FUTURE WORK, AND BROADER IMPACT
VII-A Current Limitations
While Chronos marks a significant advance in autonomous code maintenance, certain limitations remain:
  • Extreme-Scale Context Latency: Despite highly efficient memory and retrieval, performance can be constrained in ultra-large monolithic repositories during peak concurrent update hours.
  • Initial Memory Cold Start: Brand new projects (no history) or projects with very sparse documentation may exhibit reduced multi-code association richness, necessitating a brief period of active learning.
  • Non-Determinism and Overcorrection: In some edge cases, especially with noisy historical feedback, Chronos has shown a tendency to ”over-correct” or introduce minor regressions before self-stabilization.
  • Opaque Reasoning Pathways: The fully autonomous nature and integration of deep learning sometimes hinder model explainability, which may limit adoption in highly regulated or safety-critical industries.
VII-B Future Work
We are actively addressing these limitations and have several concrete directions for improvement:
  • Incremental Embedding Optimization: Research into hybrid memory systems and smarter incremental re-indexing to further minimize retrieval and context assembly latency during scale-out operations.
  • Interactive and Visual Explanations: Implementation of stepwise reasoning visualizers and debug trace replays, making Chronos’ decision process more transparent to engineers and auditors.
  • Multi-Project/Organization Federated Memory: Extending Chronos memory to operate across multiple repositories and organizational boundaries, supporting federated bug detection, code migration, and systemic risk profiling.
  • Human-in-the-Loop Collaboration: Exploring adaptive workflows where engineers and Chronos can jointly control the loop—overriding, annotating, or confirming automated fixes in sensitive code regions.
  • Security and Adversarial Robustness: Building rigorous defense mechanisms against prompt injection, memory poisoning, and adversarial debugging cases.
VII-C Deployment Architecture and Integration
@@ VII-D Broader Impact
VIII CONCLUSION
@@
Our comprehensive evaluation reveals that Chronos achieves a 65.3% success rate on real-world debugging benchmarks, representing a 6-7x improvement over state-of-the-art models including those with million-token contexts. This performance gain is attributed to three key technical contributions: (1) domain-specific pre-training on 15 million debugging instances including stack traces, fix commits, and CI/CD logs, (2) a persistent memory architecture that maintains cross-session knowledge of project-specific patterns, and (3) a hierarchical retrieval mechanism that enables effective reasoning over repository-scale contexts without computational constraints.
The implications of this work extend beyond immediate debugging applications. By demonstrating that specialized architectures and training regimes can dramatically improve performance on complex software engineering tasks, we provide evidence for the viability of task-specific language models in technical domains. Future research directions include extending this approach to other software engineering workflows, investigating transfer learning between debugging domains, and exploring human-AI collaborative debugging frameworks.

### v2

@@ VI-B Key Architectural Insights: Memory, Iteration, and Context
Bug Scenario GPT-4.1 Chronos Chronos Resolution Path
@@
TABLE XXV: Qualitative examples where Chronos successfully applies multi-code context to resolve debugging tasks beyond the reach of baseline LLMs.
VI-C Ablation Studies
@@
TABLE XXVI: Ablation study showing contribution of each component. ✓ indicates component enabled.
Model Variant Memory AGR Test Loop Fix Rate↑ Bug Loc.↑
Full Chronos 65.3% 87.2%
w/o Memory $\times$ 48.1% (-17.2) 71.5% (-15.7)
w/o AGR $\times$ 42.6% (-22.7) 63.8% (-23.4)
w/o Test Loop $\times$ 51.7% (-13.6) 79.3% (-7.9)
w/o Memory+AGR $\times$ $\times$ 31.2% (-34.1) 52.1% (-35.1)
w/o Memory+Test $\times$ $\times$ 35.8% (-29.5) 58.7% (-28.5)
w/o AGR+Test $\times$ $\times$ 33.4% (-31.9) 55.2% (-32.0)
Baseline (None) $\times$ $\times$ $\times$ 22.1% (-43.2) 41.3% (-45.9)
FullNo MCAStaticNo Orch0202020404040606060808080100100100909090494949626262555555ConfigurationDebug Success (%)
@@
VI-D Failure Mode Analysis: Understanding Chronos’s Limitations
@@
VI-D1 Common Failure Modes
@@
VI-D2 Edge Cases and Limitations
  • Extremely Large Monorepos ($>10M$ LOC): Performance degrades to 45.3% success rate due to retrieval precision issues
@@ VI-D2 Edge Cases and Limitations
Concrete Failure Case Study: Chronos struggled with a TypeScript monorepo using Lerna, dynamic imports, and cross-package configuration resolution. Despite correctly retrieving all relevant files across packages, PDM lacked adequate context to resolve type scope ambiguity when multiple packages exported identically-named interfaces. The fix attempted to import from the wrong package, causing circular dependencies. This failure mode highlights the need for semantic-aware static linking and better understanding of module resolution strategies in complex build systems. Future work should incorporate build tool semantics directly into the memory layer.
@@
TABLE XXVII: Chronos performance on challenging bug categories.
VI-E Adversarial Evaluation
To assess Chronos’s robustness against malicious inputs and poisoned training data, we conduct comprehensive adversarial testing:
VI-E1 Attack Methodology
We evaluate three categories of adversarial attacks:
1. Input Perturbation Attacks:
  • Prompt Injection: Inserting malicious instructions in bug descriptions
  • Code Obfuscation: Deliberately confusing variable names and control flow
  • Misleading Comments: Comments that contradict actual code behavior
2. Data Poisoning Attacks:
  • Backdoor Triggers: Specific code patterns that trigger incorrect fixes
  • Label Flipping: Training examples with intentionally wrong fixes
  • Gradient Attacks: Adversarial examples crafted to maximize loss
3. Retrieval Manipulation:
  • Context Stuffing: Flooding retrieval with irrelevant but similar code
  • Dependency Confusion: Creating fake dependencies to mislead AGR
  • Temporal Attacks: Manipulating commit timestamps to affect retrieval
VI-E2 Robustness Results
TABLE XXVIII: Adversarial robustness evaluation on 1,000 attack samples per category
Category Attack Type Success Detection Mitigation
Rate Rate
Input Prompt Injection 12.3% 87.7% Input sanitization
Perturbation Code Obfuscation 23.1% 76.9% AST normalization
Misleading Comments 8.7% 91.3% Code-comment align
Data Backdoor Triggers 5.2% 94.8% Anomaly detection
Poisoning Label Flipping 15.6% 84.4% Consistency checks
Gradient Attacks 19.3% 80.7% Gradient clipping
Retrieval Context Stuffing 31.4% 68.6% Relevance filtering
Manipulation Dependency Confusion 27.8% 72.2% Graph validation
Temporal Attacks 11.2% 88.8% Timestamp verify
Key Findings:
  • AGR’s graph structure provides natural defense against context stuffing (68.6% detection)
  • PDM’s pattern matching detects 94.8% of backdoor triggers through anomaly scores
  • Temporal attacks are largely ineffective due to multi-signal validation
VI-F Scalability Analysis
We evaluate Chronos’s performance across codebases ranging from 1K to 100M lines of code:
VI-F1 Performance Metrics by Scale
TABLE XXIX: Scalability analysis across different codebase sizes
Codebase Size Debug Avg Response Memory Graph Build Index
(LOC) Success Time (s) Usage (GB) Time (min) Size (GB)
1K-10K 91.2% 2.3 0.5 0.1 0.01
10K-100K 89.7% 4.7 2.1 1.2 0.08
100K-1M 87.1% 8.9 8.7 12.4 0.92
1M-10M 82.3% 18.2 31.5 87.3 9.7
10M-100M 73.4% 45.7 128.3 512.8 98.2
VI-F2 Optimization Strategies at Scale
  1. 1.
    Incremental Graph Updates: For codebases $>1M$ LOC, we implement differential graph updates that process only changed files, reducing update time by 87%.
  2. 2.
    Hierarchical Indexing: Multi-level indexes with module-level summaries enable sub-linear retrieval complexity O(log n) for repositories $>10M$ LOC.
  3. 3.
    Distributed Processing: Graph construction parallelizes across 32 cores, achieving near-linear speedup for codebases up to 50M LOC.
  4. 4.
    Memory-Mapped Storage: PDM uses memory-mapped files for patterns exceeding RAM, maintaining <100<100< 100ms access latency.
VI-F3 Detailed Memory Consumption Analysis
We profile memory usage across different components and operations:
TABLE XXX: Memory consumption breakdown by component (1M LOC codebase)
Component Peak Memory % of Total Growth Rate
AGR Graph Structure 2.8 32.2% O(n log n)
PDM Pattern Storage 1.9 21.8% O(m)
Embedding Cache 1.6 18.4% O(n)
AST Representations 1.2 13.8% O(n)
Temporal Indexes 0.7 8.0% O(n log t)
Runtime Buffers 0.5 5.8% O(1)
Total 8.7 100% -
Memory Optimization Strategies:
  • Lazy Loading: Graph nodes loaded on-demand, reducing baseline memory by 65%
  • Pattern Compression: PDM patterns compressed using AST-aware encoding (3.2x compression ratio)
  • Cache Eviction: LRU policy for embeddings with 90% hit rate at 20% memory cost
  • Incremental GC: Custom garbage collection for graph traversal reduces peak memory by 40%
VI-G Human Evaluation Study
We conducted a controlled study with N=50 professional developers to assess Chronos’s real-world effectiveness:
VI-G1 Study Design
Participants: 50 developers (5-15 years experience) from 12 companies
  • 20 backend engineers (Java/Python)
  • 15 full-stack developers (JavaScript/TypeScript)
  • 10 infrastructure engineers (Go/Rust)
  • 5 ML engineers (Python/C++)
Tasks: Each developer debugged 10 real production bugs:
  • 5 bugs with Chronos assistance
  • 5 bugs with their preferred tools (baseline)
  • Randomized order to prevent learning effects
  • Bugs selected from actual production incidents
VI-G2 Quantitative Results
TABLE XXXI: Human evaluation results (N=50 developers, 500 debugging sessions)
Metric With Chronos Baseline Improvement p-value
Fix Success Rate 87.2% 62.4% +39.7% <0.001<0.001< 0.001
Time to Fix (min) 18.3±7.2 43.7±19.8 -58.1% <0.001<0.001< 0.001
Code Quality Score 8.7/10 7.2/10 +20.8% <0.01<0.01< 0.01
Confidence Rating 8.9/10 6.8/10 +30.9% <0.001<0.001< 0.001
Would Use Again 92% - - -
VI-G3 Qualitative Feedback
Positive Themes:
  • ”AGR found cross-file dependencies I would have missed” (31/50 developers)
  • ”PDM patterns saved hours on recurring issues” (27/50)
  • ”Explanations helped me understand the fix” (42/50)
Areas for Improvement:
  • ”Needs better IDE integration” (18/50)
  • ”Sometimes retrieves too much context” (12/50)
  • ”UI lag on very large codebases” (8/50)
VI-H Fine-grained Ablations
We conduct detailed ablations on individual AGR components to understand their contributions:
VI-H1 Component-wise Ablation Results
TABLE XXXII: Fine-grained ablation study on AGR components (MRR benchmark)
Category Configuration Debug Prec@3 Hops
Baseline Full AGR System 87.1% 92.8% 2.3
Graph Remove AST edges 71.2% 84.1% 3.1
Construction Remove log flow edges 78.9% 87.3% 2.7
Remove test trace edges 73.4% 81.2% 3.8
Remove temporal weights 82.3% 89.7% 2.5
Traversal Fixed k=3 (no adaptive) 69.8% 78.4% 3.0
Strategy No confidence weighting 74.1% 83.2% 4.2
BFS instead of guided 61.3% 71.8% 5.7
No backtracking 79.2% 88.1% 2.8
Memory No PDM patterns 72.8% 85.3% 3.2
Integration No pattern ranking 81.2% 90.1% 2.6
No temporal decay 83.7% 91.2% 2.4
Critical Findings:
  1. 1.
    AST edges are most critical: Removing them causes 18.3% performance drop
  2. 2.
    Adaptive depth is essential: Fixed k=3 reduces success by 19.9%
  3. 3.
    PDM integration provides 16.4% boost: Pattern memory significantly improves debugging
  4. 4.
    Guided traversal beats BFS by 29.6%: Intelligence in path selection matters
VII DESIGN RATIONALE AND THEORETICAL MOTIVATION FOR AGR
The Adaptive Graph-Guided Retrieval (AGR) system addresses three fundamental limitations of vector-based retrieval in debugging contexts. First, real codebases exhibit structural dependencies that span multiple files, a bug in AuthService.java may stem from a configuration change in config.yml that affects DatabasePool.java, which then impacts authentication. Traditional top-k vector search fails here because these files share minimal textual similarity. AGR’s multi-hop traversal follows actual code dependencies (imports, function calls, data flow), recovering the complete causal chain regardless of embedding distances. Second, local neighborhood propagation in AGR exploits the insight that code proximity (in the dependency graph) correlates strongly with debugging relevance. When a test fails in PaymentTest.java, the bug is exponentially more likely to be within 1-3 hops in the dependency graph than in a random file with high textual similarity. This locality principle, combined with confidence-weighted traversal, enables AGR to achieve 89% precision where vector search achieves only 31%. Third, temporal anchoring prevents the retrieval of stale code patterns in actively maintained repositories. By weighting edges based on commit recency and co-modification frequency, AGR naturally prioritizes current architectural patterns over deprecated ones, reducing false positives by 67% in repositories with >1000>1000> 1000 commits/month.
Bug: NPE in ExportExportService.javaExportUtils.javaExportConfig.java×$\times$ Misses Authsim=0.89sim=0.72sim=0.68Vector RAG: Top-3 SimilarBug: NPE in ExportExportService.javaAuthService.javaTokenCache.java✓ Found Root Causehop=0callsusesrecentcommitAGR: Dependency Traversal3x better
Figure 19: AGR vs Vector RAG: While vector search retrieves textually similar files that miss the root cause, AGR follows actual code dependencies and temporal signals to locate the authentication bug causing the export failure.
The superiority of AGR demonstrated in Figure 19 is not merely empirical, it stems from fundamental theoretical properties that we now formalize.
VIII THEORETICAL GUARANTEES
Having demonstrated AGR’s practical superiority through extensive evaluation, we now provide theoretical foundations that explain why these improvements are fundamental rather than incidental.
VIII-A Why AGR Converges Efficiently
[…]
Convergence Guarantee: Given a code graph with maximum degree $d$, AGR converges to the optimal debugging context within $k$ iterations (typically k=5) with probability exceeding 98%.
Practical Implications:
  • Bounded search: AGR examines at most $O(d^{k})$ nodes, preventing exponential blowup
  • Early termination: 73% of bugs are resolved within 3 hops
  • Predictable performance: Response time scales logarithmically with codebase size
VIII-B Fix Correctness with PDM
The Persistent Debug Memory provides statistical guarantees on fix quality:
Pattern Matching Accuracy: With $m$ stored bug patterns, the probability of generating a correct fix increases as:
$P(\text{correct fix})\geq 1-e^{-m/1000}$
[…]
VIII-C Why This Matters for Production Debugging
Unlike theoretical debugging models, AGR’s guarantees translate directly to production benefits:
TABLE XXXIII: Theoretical guarantees vs. real-world impact
Theoretical Property Guarantee Production Impact
Convergence time O(k log d) <10<10< 10s for 1M LOC
[…]
Memory usage O(n log n) Fits in 32GB RAM
[…]
IX FAILURE ANALYSIS: DEEP DIVE INTO CHALLENGING CASES
While our theoretical guarantees and empirical results demonstrate Chronos’s effectiveness, understanding failure modes is crucial for both users and future improvements. We conduct systematic failure analysis on 2,500 debugging sessions where Chronos failed to produce correct fixes:
IX-A Failure Taxonomy
TABLE XXXIV: Detailed failure analysis across 2,500 failed debugging sessions
Failure Category Subcategory Count % of Failures Root Cause
Retrieval Failures
(42.3% of total)
Missing Context 423 16.9% AGR depth limit reached
Wrong Context 312 12.5% Semantic aliasing confusion
Stale Context 323 12.9% Outdated cached patterns
Understanding Failures
(31.2% of total)
Complex Logic 287 11.5% Multi-step reasoning limit
Domain Specific 234 9.4% Missing specialized knowledge
Implicit Behavior 256 10.2% Undocumented assumptions
Generation Failures
(26.5% of total)
Incorrect Fix 298 11.9% Wrong root cause identified
Partial Fix 189 7.6% Incomplete solution
Breaking Changes 176 7.0% Side effects not considered
IX-B Representative Failure Cases
IX-B1 Case 1: Distributed Consensus Bug
Bug Description: Raft consensus implementation fails during leader election with 5+ nodes
Failure Mode: Chronos retrieved all relevant consensus code but failed to identify the subtle race condition in vote counting across network partitions.
// Chronos attempted fix (incorrect)
if (voteCount > clusterSize / 2) {
becomeLeader(); // Missing: partition check
// Correct fix required understanding of:
// 1. Network partition detection
// 2. Split-brain prevention
// 3. Distributed state consistency
Listing 1: Distributed consensus bug: Chronos’s incorrect fix
Root Cause: Lacks theoretical understanding of distributed systems safety properties and FLP impossibility theorem implications.
IX-B2 Case 2: Memory Corruption in C++ Template Metaprogramming
Bug Description: Segfault in recursive template instantiation with variadic parameters
Failure Mode: AGR correctly identified template definitions but PDM had no patterns for template instantiation depth issues.
// Complex template causing stack overflow
template<typename... Args>
struct TypeList {
template<template<typename...> class F>
using apply = F<Args...>;
// Chronos missed: recursive instantiation depth
// Required: SFINAE and constexpr evaluation
Listing 2: C++ template metaprogramming failure
Root Cause: Template metaprogramming requires compile-time reasoning that exceeds current model capabilities.
IX-B3 Case 3: Machine Learning Pipeline Data Leakage
Bug Description: Validation accuracy drops 40% in production despite 95% test accuracy
Failure Mode: Chronos identified standard preprocessing steps but missed subtle data leakage through feature normalization before train/test split.
Root Cause: Requires understanding of statistical independence and temporal data dependencies in ML pipelines.
IX-C Failure Pattern Analysis
11122233344455566677788899910101005050501001001002.35.18.715.223.638.952.171.384.793.2Bug Complexity (1-10 scale)Failure Rate (%)
Figure 20: Failure rate increases exponentially with bug complexity. The system struggles with highly complex bugs (complexity 8+) that require cross-module understanding.
IX-D Mitigation Strategies and Future Directions
1. Enhanced Retrieval for Edge Cases:
  • Implement fallback to exhaustive search when confidence <0.3<0.3< 0.3
  • Add cross-repository pattern matching for rare bugs
  • Develop specialized retrievers for distributed/concurrent code
2. Improved Understanding through Neuro-Symbolic Integration:
  • Integrate theorem provers for correctness verification
  • Add symbolic execution for path exploration
  • Incorporate domain-specific reasoning modules
3. Generation Safety Mechanisms:
  • Mandatory test generation before fix application
  • Rollback mechanisms for breaking changes
  • Human-in-the-loop validation for critical systems
The Bottom Line Despite these edge cases, Chronos achieves 87.1% debugging success compared to 22.9% for the next best system. Even in failure cases, Chronos provides actionable insights that accelerate manual debugging by 2.3x on average.
X LIMITATIONS AND FUTURE WORK
X-A Technical Limitations
While Chronos advances autonomous debugging capabilities, several technical constraints remain:
TABLE XXXV: Summary of Technical Limitations
Limitation Impact Affected Scenarios Mitigation
Extreme-scale latency 5s+ retrieval time Repos $>10M$ LOC Parallel expansion
Memory cold start 23% lower success New projects $<1K$ commits Transfer learning
Non-determinism 18.3% variance Distributed systems Deterministic replay
Reasoning opacity 10GB traces/bug All debugging sessions Selective logging
Memory coherence 2.1% conflicts Concurrent instances Eventual consistency
Dynamic languages 41.2% accuracy Python/Ruby/JS Runtime instrumentation
External dependencies 31% lower success API/DB bugs Mock generation
Key Technical Constraints:
  • Extreme-Scale Context Latency: O(k²n) complexity causes 5s+ retrieval times in 10M+ LOC repositories.
  • Memory Cold Start: 23% lower success on projects with $<1K$ commits until memory builds over 2-3 weeks.
  • Non-Determinism: 18.3% variance in distributed systems due to timing and resource contention.
  • Reasoning Transparency: 10GB trace data per bug makes full interpretability infeasible.
  • Dynamic Languages: 41.2% accuracy on Python/Ruby/JS due to runtime metaprogramming.
  • External Dependencies: 31% lower success on API/database bugs without visibility into external states.
X-A1 Failure Mode Analysis
TABLE XXXVI: Common failure modes, frequencies, and mitigation strategies.
Failure Mode Root Cause Frequency Mitigation Strategy
Incomplete fixes Insufficient test coverage 12.3% Automated test generation
Over-engineering Historical pattern bias 8.7% Confidence thresholding
Context overflow Repository complexity 6.2% Hierarchical retrieval
False positives Ambiguous error messages 4.8% Multi-source validation
Regression introduction Side effect blindness 3.1% Impact analysis expansion
X-B Research Directions: Extending the Debugging Paradigm
Key research directions:
  • Algorithmic Optimization: Sub-quadratic retrieval for 10M+ LOC repositories
  • Visual Understanding: Screenshot analysis for UI/UX debugging
  • Federated Learning: Cross-organization bug patterns with privacy
  • Human-AI Collaboration: Interactive debugging with feedback loops
  • Security Hardening: Defense against adversarial attacks
X-C Deployment Architecture and Integration
@@ X-D Broader Impact
DebugSuccessRoot CauseAccuracyRetrievalPrecisionMulti-FileUnderstandingFixQuality020202040404060606080808010010010071.289.283.578.9Evaluation MetricPerformance (%)ChronosClaude 4 OpusGPT-4.1Gemini 2.0 ProDeepSeek V3
Figure 21: Comprehensive performance comparison across key debugging metrics. Chronos consistently outperforms state-of-the-art models by 3-5x across all evaluation dimensions.
TABLE XXXVII: End-to-end debugging pipeline efficiency metrics
System Avg. Time to Fix Iterations Token Usage Cost per Bug Success Rate
Chronos 4.2 min 2.2 31.2K \$0.18 65.3%
Claude 4 Opus + RAG 18.7 min 5.8 142.3K \$2.84 14.2%
GPT-4.1 + LangChain 21.3 min 6.2 156.7K \$3.13 13.8%
Gemini 2.0 Pro 19.5 min 5.5 134.8K \$1.35 13.4%
Human Developer 35.8 min 3.4 N/A \$29.83 87.2%
Based on average developer hourly rate of \$50
XI CONCLUSION
@@
Our comprehensive evaluation reveals that Chronos achieves a 65.3% success rate on real-world debugging benchmarks, representing a 4-5x improvement over the best 2025 models including Claude 4 Opus (14.2%), GPT-4.1 (13.8%), and Gemini 2.0 Pro (13.4%). This performance gain persists despite these models achieving state-of-the-art results on code generation tasks (Claude 4 Opus: 72.5% on SWE-bench). The disparity confirms our hypothesis: debugging is fundamentally different from code generation and requires specialized architectures.
Key technical contributions enabling this breakthrough include: (1) domain-specific pre-training on 15 million debugging instances including stack traces, fix commits, and CI/CD logs, (2) Adaptive Graph-Guided Retrieval (AGR) that outperforms advanced RAG techniques like HyDE, Self-RAG, and FLARE by 2-3x on debugging tasks, (3) a persistent memory architecture that maintains cross-session knowledge, a capability absent in modern IDEs like Cursor and Windsurf, and (4) an autonomous debugging loop with iterative refinement based on test execution feedback.
The implications of this work extend beyond immediate debugging applications. By demonstrating that specialized architectures and training regimes can dramatically improve performance on complex software engineering tasks, we provide evidence for the viability of task-specific language models in technical domains. While Chronos is built for debugging, its architecture naturally extends to long-term codebase governance , enabling intelligent memory of design decisions, bug patterns, and recovery heuristics. Future research directions include extending this approach to other software engineering workflows, investigating transfer learning between debugging domains, and exploring human-AI collaborative debugging frameworks.